(p/4)-(3p/8)=4 solve the equation.
Found 2 solutions by nyc_function, stanbon:
Answer by nyc_function(2741)

LCD = 8
(p/4)* 8 -(3p/8)* 8 = 4*8
2p - 3p = 32
-p = 32
p = 32/-1
p = -32

Answer by stanbon(75887)

(p/4)-(3p/8)=4
==================
When you have an equation it is easy to
get rid of the fractions.
---------------
Multiply thru the equatiion by 8 to get:
2p - 3p = 32
-p = 32
p = -32
==================
Cheers,
Stan H.
==================